OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!! What is the exact area of a circle with a diameter of 25 m?

OpenStudy (campbell_st):

find the radius then use \[A = \pi \times r^2\] where r = radius

OpenStudy (campbell_st):

the alternative us to let r = 25/2 then its \[A = \pi \times (\frac{25}{2})^2\]

OpenStudy (anonymous):

formula for area is pi*d^2/4 pi is 3.14 d=25 so area is( 3.14*25*25)all over4 =490.87

OpenStudy (campbell_st):

but that's not exact.... make it \[A = \pi \times \frac{25^2}{2^2}\] and leave the answer as a fraction

OpenStudy (anonymous):

yeah i agree with u i dont think it is 490 either i think it is 49.73 @campbell_st

OpenStudy (campbell_st):

well what is 25^2 = ?

OpenStudy (anonymous):

50 @campbell_st

OpenStudy (campbell_st):

no you need 25 x 25 =

OpenStudy (anonymous):

i mean 625 @campbell_st

OpenStudy (campbell_st):

great so the exact area is \[A = \frac{625 \times \pi}{4}\] all you need to do is remove the multiplication sign from the numerator to get and exact value

OpenStudy (anonymous):

thanks :) @campbell_st

OpenStudy (radar):

You left off something that some teachers would take off some. The answer did not include the units for area, in this case sq meters.